OpenStudy (anonymous):

George Yilling bought a share of stock at 41 3/8 on Monday. On Tuesday it cost 43 1/8 per share. Find the net change from Monday to Tuesday.

OpenStudy (paki):

do you know, how to write 41(3/8) in simple from.....

OpenStudy (anonymous):

no i don't.

OpenStudy (paki):

ok... 41(3/8) will be equal to 41.375....

OpenStudy (paki):

and 43 (1/8) will be equal to 41.125.... got this...?

OpenStudy (anonymous):

um okay.

OpenStudy (anonymous):

@paki isnt 43(1/8) = 43.125

OpenStudy (paki):

now "net change" can be find out by (41.375/43.125)*100 = 95.94% ...

OpenStudy (anonymous):

I got it! 43.125-41.125=1.75.

OpenStudy (anonymous):

i feel @epgdanii is right

OpenStudy (paki):

@Rohith15 yup it is 43.125... typing mistake....

OpenStudy (paki):

yup and 1.75 is 95.94% of the 43..125

OpenStudy (anonymous):

Thank you guys so much:)!

OpenStudy (paki):

pleasure :-)